solve for a: -(a+8)=5(a-1)+6

solve for a: (a+8)=5(a-9)+17

Answer by rfer(16322)

-a-8=5a-5+6
-6a=9
a=-1.5
-----------
a+8=5a-45+17
-4a=-36
a=9